UROP Proposal

Chris Shabsin

February 17, 1993

In my UROP for DCNS this term, I will be part of a team working to implement SLIP for TechMail-S. In this process, I’ll learn much about how SLIP in general works, as well as learning even more about Windows programming and I’ll gain some insight into how the TCP/IP kernel for DOS machines works.

I’ll also be writing rough drafts of a few documents about various subjects based on my previous experiences, including DCNS’s options with OS/2 and hints and notes on the development of an installation program for Lan Workplace for DOS.

Another possible project for this term will be an authenticated LPR client for DOS. This will allow people with PC’s on the net to print to athena’s printers with the standard quota system.

Projects will be added and deleted from this list as the term progresses.